# Differential Privacy for Transformer Embeddings with Nonparametric Variational Information Bottleneck

### v1

@@ 2.3 Related Work on Private Text Embeddings
The problem of sharing privacy-preserving text embeddings under local differential privacy (LDP) has received increasing attention, confirming the relevance of our problem space. Recent notable efforts include (du2023sanitizing; meehan2022sentence). (du2023sanitizing) focus on applying noise directly to sentence embeddings using DP mechanisms like Gaussian or Laplace noise, along with considering label privacy. While effective in providing LDP guarantees, its noise calibration is primarily driven by privacy parameters, in a task-agnostic manner, without explicit, learned consideration for the nuances of downstream task utility. Similarly, (meehan2022sentence) addresses privacy for document embeddings by generating differentially private sentence embeddings. This approach involves methods that introduce noise uniformly or based on sensitivity metrics, also operating under the LDP setting for textual representations.
These pieces of work make important contributions that operate in the same relevant setting as our work, demonstrating the feasibility and demand for LDP for textual data. Our primary novelty, which fundamentally distinguishes our method, is the task-aware calibration of noise via the Nonparametric Variational Information Bottleneck (NVIB). Unlike this prior work, which employs task-agnostic noise mechanisms that apply noise uniformly or based on general sensitivity to individual vectors, NVDP integrates the privacy mechanism directly into the learning process of the transformer. This allows the noise level to be dynamically adjusted based on the specific utility requirements of the downstream task, optimizing the privacy-utility trade-off over the whole multi-vector document embedding. This input-dependent learned calibration ensures that sensitive information is protected adequately, while critical task-relevant features are preserved more effectively, a capability not explicitly demonstrated or achieved by the task-agnostic approaches in (du2023sanitizing; meehan2022sentence).
@@ 3 NVDP: Nonparametric Variational Differential Privacy
Our goal is to share transformer embeddings in a privacy-preserving manner, but still share embeddings which are useful. Intuitively, RDP measures the degree of privacy by measuring how much information a sampled embedding is expected to convey (with more weight put on the worst case samples). If we know what type of task we intend the embeddings to be used for, then we can measure the degree of utility of an embedding by measuring the accuracy of a model trained on such a task. Reducing the total amount of information in a latent representation while retaining enough information to perform a task is precisely the objective of a variational information bottleneck (VIB) regulariser. VIB trains a posterior distribution over embeddings, a sample from this posterior provides limited information, and this information is optimised to perform the downstream task. More precisely, the KL-divergence used in VIB is exactly equivalent to the Renyi-divergence used in RDP as $\lambda$ approaches $1$ (at which point all samples get equal weight, including the worst case).
We apply this general idea to the case of transformer embeddings. This is complicated by the fact that transformer embeddings consist of a sequence of vectors, so we use NVIB to provide a holistic view of the amount of information in these variable-length sequences. NVIB is used as the regulariser to calibrate the noise, and RDP is used as the privacy measure to evaluate the noise. Thus, we propose nonparametric variational differential privacy (NVDP) as a method to learn posterior distributions over embeddings which provide good RDP privacy and are calibrated to a given utility objective and the empirical data distribution.
@@ 3.1 NVDP Architecture
The architecture of our proposed NVDP model is illustrated in Figure 1. NVDP builds on top of a pretrained transformer encoder [^2] We use BERT-base-uncased model to generate embeddings., which generates multi-vector transformer embeddings. These embeddings are processed by a single layer Transformer that incorporates our modified NVIB mechanism, projecting the embeddings into a mean, variance and pseudo-count for each input vector, which together determine the posterior distribution over sequences of weighted vectors. Two key modifications are made to ensure privacy. First, we sample from this posterior distribution during both training and testing to generate a noisy, sanitized version of the embedding. This stochastic bottleneck is the core of our privacy mechanism. Second, we process these noisy embeddings with a Denoising Multi-Head Attention (MHA), but we remove the standard residual skip connection that would typically wrap this block. This architectural change is critical, as it prevents any un-sanitized information from the original embedding $x$ from leaking past the noisy latent representation and into the final output. This ensures that all shared information is passed exclusively through the privacy-preserving bottleneck.
Figure 1: An NVDP model. The model projects an input embedding into the parameters of a posterior distribution. A key modification for privacy is that we sample from this latent distribution during both training and testing. The resulting noisy representation is processed by a Denoising MHA layer. To enforce the privacy bottleneck, the standard residual skip connection around the MHA is removed, preventing any un-sanitized information from bypassing the bottleneck.
@@ 3.2 Measuring Privacy
In our local differential privacy setting, we want to make sure that any sample $\bm{S}$ from the embedding $Q(\bm{S})$ of a text $x$, could also have come from a different embedding $Q^{\prime}(\bm{S})$ of some different text $\bm{x}$$^{\prime}$. We measure this basic privacy criteria using the Rényi Divergence $D_{\lambda}(Q||Q^{\prime})$, defined in Definition 2.1, to quantify the difference between these two sampling distributions. In this work, we measure privacy from two complementary perspectives, which differ in how we aggregate across the possible alternative inputs $\bm{x}$$^{\prime}$. For both these measures, we report the worst case across the given input $\bm{x}$} (the average case is reported in Table 2 in Appendix A).
The first approach measures the worst case also across all alternative inputs $\bm{x}$$^{\prime}$, as in standard RDP. We do not assume any specific notion of adjacency between examples. In our experiments, we report the maximum Renyi divergence over all input pairs as the RDP measure (the average case is reported in Table 2 in Appendix A ).
@@
To calculate the Rényi divergence (RD) between the posterior embedding $Q(\bm{S})$ for input $\bm{x}$} and an alternative embedding $Q^{\prime}(\bm{S})$ for input $\bm{x}$$^{\prime}$, we first specify a sampling procedure which approximates sampling from a Dirichlet Process, and then calculate the RD for that sampling procedure applied to $Q$ and $Q^{\prime}$. This is an upper bound on the RD between the two Dirichlet Processes, since a sequence of weighted vectors is more specific than a mixture distribution, but it is appropriate for our privacy measure because it is this sequence which is actually shared. So the RD between sampling two different finite sequences of weighted vectors from two different inputs is an accurate assessment of how easily an observer can distinguish between samples from one text’s embedding and samples from the other text’s embedding, providing a measure of privacy.
@@ 3.3 Rényi Divergence between two Sampling Distributions
Since the theory of NVIB assumes that the posterior distribution which embeds the information about the input data is a Dirichlet Process, $Q=DP(\alpha^{q}_{0},G^{q}_{0})$, we want a sampling procedure which generates finite sequences of weighted vectors which approximate sampling from this DP. For $DP(\alpha^{q}_{0},G^{q}_{0})$, defined in Equation 4, weights can be sampled using a stick breaking process parameterised by $\alpha^{q}_{0}$, and vectors can be sampled independently by first sampling a component $i$ from the mixture distribution $G^{q}_{0}$ and then sampling a vector from that component’s distribution $G^{q}_{i}=\mathcal{N}(\mu^{q}_{i},(\sigma^{q}_{i})^{2})$. But sampling from the discrete list of components $i$ makes it difficult for NVIB to do backprop through the sampling step, so instead hendersonfehr2023 propose a sampling method which first samples the total weight assigned to all the vectors from a given component $i$, and then samples weighted vectors independently from the component’s distribution $G^{q}_{i}$. They prove that the DP distribution in Equation 4 is equivalent to such a factorized distribution as follows:
$F$~~~ $=~~~\sum_{i=1}^{n+1}~{\rho}_{i}{F}_{i}$ $\bm{\rho}$~~~ $\sim~~~\operatorname*{Dir}({\alpha}^{q}_{1},\ldots,{\alpha}^{q}_{n+1})$ ${F}_{i}$~~~ $=~~~DP(G_{i}^{q},\alpha_{i}^{q})\vskip-4.30554pt$ (6)
@@
In particular, hendersonfehr2023 show that sampling the total weights $\bm{\rho_{i}}$ across $i$ can be done by sampling from a Dirichlet distribution parameterised by $\bm{\alpha^{q}}$, and the distribution of weighted vectors sampled from each component is another Dirichlet process. To generate a sample with a finite number of vectors, they assume a given bound $\kappa_{i}$ on the number of vectors sampled from each component $G^{q}_{i}$, and sample their weights with a symetric Dirichlet distribution $\operatorname*{Dir}(\frac{{\alpha}^{q}_{i}}{\kappa_{i}},\overset{\kappa_{i}}{\ldots},\frac{{\alpha}^{q}_{i}}{\kappa_{i}})$. In our experiments, one vector is sampled from each component, so $\kappa_{i}=1$.
@@
The different weighted vectors sampled from a Dirichlet Process specify a mixture distribution, which is permutation invariant. However, when a sample is generated with our sampling procedure, the weighted vectors are output in a specific order. We could have a better approximation to a Dirichlet Process by randomly permuting the weighted vectors, but for simplicity we output them in the order of the tokens in the sentence. This simplifies the computation of the RD between two samples because it implies that sampled vectors for two different inputs can be aligned by their token position, which gives us an upper bound on the Dirichlet Process case, since the ordered list is more informative. [^3] To handle the case where the two examples have different numbers of tokens, we pad the inputs sequences so that they all have the same length, and then assume that pad tokens have parameters $\mu_{i}{=}\mathbf{0},~\sigma_{i}{=}\mathbf{1},~\alpha_{i}{=}\mathbf{0}$. Alternatively, we could have defined adjacent inputs as all having the same length, but this approach is more appropriate for the BDP measure and still gives a meaningful privacy measure. We leave better bounds on the RD between samples from Dirichlet Processes to future work. This gives us the following formula for calculating the Rényi divergence between samples from two different input embeddings.
@@
$+~\sum_{i=1}^{n+1}\kappa_{i}\Biggl(\frac{\lambda}{2}\Biggl\lVert\frac{\bm{\mu^{q}_{i}}-\bm{\mu^{q^{\prime}}_{i}}}{\bm{\sigma^{\prime}_{i}}}\Biggr\rVert^{2}~+~\frac{1}{1-\lambda}\bm{1}\left(\log\frac{\bm{\sigma^{\prime}_{i}}}{(\bm{\sigma_{0}^{p}})^{(1-\lambda)}(\bm{\sigma^{q}_{i}})^{\lambda}}\right)\Biggr)\vskip-2.15277pt$
@@ 4 Experiments
We empirically evaluate NVDP on NLP tasks from the GLUE benchmark (wang2018glue) in terms of both accuracy and privacy. This is a commonly-used benchmark for evaluating performance on a variety of natural language understanding tasks.
@@ 4.1 Experimental Setup
We evaluate the performance of NVDP on the General Language Understanding Evaluation (GLUE) (wang2018glue) benchmark, which is a collection of different natural language understanding tasks including similarity and paraphrasing tasks, text classification, and natural language inference (NLI). For NLI, we experiment on QNLI (rajpurkar-etal-2016-squad) and RTE (dagan2005pascal). For paraphrase detection, we evaluate on MRPC (dolan-brockett-2005-automatically), STS-B (cer-etal-2017-semeval), and QQP (chen2018quora). For text classification, we evaluate on SST-2 (socher-etal-2013-recursive).
@@
We use BERTBase, which is $12$ layers and 110110M parameters. We use the following hyper-parameters for fine-tuning BERT: a sequence length of $512$, a train batch size of $64$ and evaluation batch size of $8$. We use the stable variant of the Adam optimizer (zhang2020revisiting; mosbach2020stability) with the learning rate of 2e72e{-}7 through all experiments. We use a warm-up step of $0.2$.
@@
  • Baseline (Base): we use vanilla BERTBase model without any regularizaton technique.
@@
We consider several simplified versions of our proposed model and other privacy-preserving baselines to comprehensively evaluate the advantages of NVDP. These ablations and baselines allow us to isolate the impact of NVIB’s nonparametric regularization and task-aware noise calibration. One ablation (VTDP) removes the use of Bayesian nonparametrics for regularisation, and two ablations provide reference models which can only share single-vector embeddings instead of multi-vector transformer embeddings.
  • Variational Transformer Differential Privacy (VTDP): We consider a simplified version of our proposed model where the noise introduced with NVIB is replaced with noise introduced by vector-space VIB applied to each token vector independently, which we call the variational transformer differential privacy (VTDP) model. This provides a strong reference model which is the same as our NVDP model but without the holistic regularisation of the entire sequence of vectors provided by NVIB. We utilize BERTBase with an additional Variational Information Bottleneck (VIB) layer to implement the information bottleneck principle outlined by mahabadi2021variational. The VIB layer encodes input representations in a compressed latent space, leveraging VIB to learn a stochastic latent variable. This variable captures task-relevant information while filtering out irrelevant features, enhancing generalization. This integration enables effective fine-tuning, particularly in low-resource scenarios, by prioritizing essential information and mitigating overfitting risks. The compressed latent representation is compared to a Gaussian prior, ensuring a structured and regularized latent space. The Rényi Differential Privacy (RDP) guarantee, defined between Gaussian distributions, follows the formulation in Equation 10 of van2014renyi:
@@
    $=$ $\frac{\lambda}{2}\Biggl\lVert\frac{\bm{\mu^{q}_{i}}-\bm{\mu^{q^{\prime}}_{0}}}{\bm{\sigma^{\prime}_{i}}}\Biggr\rVert^{2}+\frac{1}{1-\lambda}\bm{1}\left(\log\frac{\bm{\sigma_{i}^{\prime}}}{(\bm{\sigma_{0}^{q^{\prime}}})^{(1-\lambda)}(\bm{\sigma^{q}_{i}})^{\lambda}}\right)$
@@
    where the noise scale $\sigma$ is fixed and shared across all dimensions and inputs (we use $\sigma=0.55$). In contrast to our NVDP model, which learns a data-adaptive and task-aware noise structure via NVIB, VIB-fixed applies a uniform and input-independent perturbation to a single vector. This serves as a transparent benchmark closely aligned with classical Gaussian Differential Privacy and admits closed-form Rényi divergence computations. For two adjacent inputs $x$ and $x^{\prime}$, the order-$\lambda$ Rényi divergence between their noisy distributions is given by mironov2017renyi:
@@
    This provides an analytically grounded reference point for comparison with NVDP, allowing us to report summary statistics of these pairwise Rényi divergences across all evaluated embeddings.
@@
    Thus, VIB-learned samples from a Gaussian distribution with mean $\mu(x)$ and diagonal covariance $\mathrm{diag}(\bm{\sigma}^{2})$, where the noise magnitudes are learned globally during training but remain input-independent. This parallels the VIB-fixed mechanism while enabling dimension-wise adaptive perturbation. For two adjacent inputs $x$ and $x^{\prime}$, the order-$\lambda$ Rényi divergence between their privatized distributions is mironov2017renyi:
    $D_{\lambda}\!\left(\mathcal{N}(\mu(x),\mathrm{diag}(\bm{\sigma}^{2}))\,\big\|\,\mathcal{N}(\mu(x^{\prime}),\mathrm{diag}(\bm{\sigma}^{2}))\right)=\frac{\lambda}{2}\left\lVert\frac{\mu(x)-\mu(x^{\prime})}{\bm{\sigma}}\right\rVert_{2}^{2}$ (10)

### v2

@@ 2.3 Related Work on Private Text Embeddings
The problem of sharing privacy-preserving text embeddings under LDP has received increasing attention, confirming the relevance of our problem space. Recent notable efforts include Du et al. (2023); Meehan et al. (2022). Du et al. (2023) focuses on applying noise directly to sentence embeddings using DP mechanisms like Gaussian or Laplace noise, along with considering label privacy. While effective in providing LDP guarantees, its noise calibration is primarily driven by privacy parameters, in a task-agnostic manner, without explicit, learned consideration for the nuances of downstream task utility. Similarly, Meehan et al. (2022) addresses privacy for document embeddings by generating differentially private sentence embeddings. This approach involves methods that introduce noise uniformly or based on sensitivity metrics, also operating under the LDP setting for textual representations.
These pieces of work make important contributions that operate in the same relevant setting as our work, demonstrating the feasibility and demand for LDP for textual data. Our primary novelty, which fundamentally distinguishes our method, is the task-aware calibration of noise via NVIB. Unlike this prior work, which employs task-agnostic noise mechanisms that apply noise uniformly or based on general sensitivity to individual vectors, NVDP integrates the privacy mechanism directly into the learning process of the transformer. This allows the noise level to be dynamically adjusted based on the specific utility requirements of the downstream task, optimizing the privacy-utility trade-off over the whole multivector document embedding. This input-dependent learned calibration ensures that sensitive information is protected adequately, while critical task-relevant features are preserved more effectively, a capability not explicitly demonstrated or achieved by the task-agnostic approaches in Du et al. (2023); Meehan et al. (2022).
@@ 3 NVDP: Nonparametric Variational Differential Privacy
Our goal is to share transformer embeddings in a privacy-preserving manner, but still share embeddings which are useful. Intuitively, RDP measures the degree of privacy by measuring how much information a sampled embedding is expected to convey (with more weight put on the worst-case samples). If we know what type of task we intend the embeddings to be used for, then we can measure the degree of utility of an embedding by measuring the accuracy of a model trained on such a task. Reducing the total amount of information in a latent representation while retaining enough information to perform a task is precisely the objective of a VIB regularizer. VIB trains a posterior distribution over embeddings, a sample from this posterior provides limited information, and this information is optimized to perform the downstream task. More precisely, the KL divergence used in VIB is exactly equivalent to the RD used in RDP as $\lambda$ approaches $1$ (at which point all samples get equal weight, including the worst-case).
We apply this general idea to the case of transformer embeddings. This is complicated by the fact that transformer embeddings consist of a sequence of vectors, so we use NVIB to provide a holistic view of the amount of information in these variable-length sequences. NVIB is used as the regularizer to calibrate the noise, and RDP is used as the privacy measure to evaluate the noise. Thus, we propose NVDP as a method to learn posterior distributions over embeddings which provide good RDP privacy and are calibrated to a given utility objective and the empirical data distribution.
@@ 3.1 NVDP Architecture
The architecture of our proposed NVDP model is illustrated in Figure 1. NVDP builds on top of a pretrained transformer encoder [^2] We use BERT-base-uncased model to generate embeddings., which generates multivector transformer embeddings. These embeddings are processed by a single layer Transformer that incorporates our modified NVIB mechanism, projecting the embeddings into a mean, variance and pseudo-count for each input vector, which together determine the posterior distribution over sequences of weighted vectors. Two key modifications are made to ensure privacy. First, we sample from this posterior distribution during both training and testing to generate a noisy, sanitized version of the embedding. This stochastic bottleneck is the core of our privacy mechanism. Second, we process these noisy embeddings with a Denoising Multi-Head Attention (MHA), but we remove the standard residual skip connection that would typically wrap this block. This architectural change is critical, as it prevents any unsanitized information from the original embedding $x$ from leaking past the noisy latent representation and into the final output. This ensures that all shared information is passed exclusively through the privacy-preserving bottleneck.
Figure 1: An NVDP model. The model projects an input embedding into the parameters of a posterior distribution. A key modification for privacy is that we sample from this latent distribution during both training and testing. The resulting noisy representation is processed by a Denoising MHA layer. To enforce the privacy bottleneck, the standard residual skip connection around the MHA is removed, preventing any unsanitized information from bypassing the bottleneck.
@@ 3.2 Measuring Privacy
In our LDP setting, we want to make sure that any sample $\bm{S}$ from the embedding $Q(\bm{S})$ of a text $x$, could also have come from a different embedding $Q^{\prime}(\bm{S})$ of some different text $\bm{x}$$^{\prime}$. We measure this basic privacy criteria using the RD $D_{\lambda}(Q||Q^{\prime})$, defined in Definition 2.1, to quantify the difference between these two sampling distributions. In this work, we measure privacy from two complementary perspectives, which differ in how we aggregate across the possible alternative inputs $\bm{x}$$^{\prime}$. For both these measures, we report the worst case across the given input $\bm{x}$} (the average case is reported in Table 2 in Appendix A).
The first approach measures the worst-case also across all alternative inputs $\bm{x}$$^{\prime}$, as in standard RDP. We do not assume any specific notion of adjacency between examples. In our experiments, we report the maximum RD over all input pairs as the RDP measure (the average-case is reported in Table 2 in Appendix A ).
@@
To calculate the RD between the posterior embedding $Q(\bm{S})$ for input $\bm{x}$} and an alternative embedding $Q^{\prime}(\bm{S})$ for input $\bm{x}$$^{\prime}$, we first specify a sampling procedure which approximates sampling from a Dirichlet Process, and then calculate the RD for that sampling procedure applied to $Q$ and $Q^{\prime}$. This is an upper bound on the RD between the two Dirichlet Processes, since a sequence of weighted vectors is more specific than a mixture distribution, but it is appropriate for our privacy measure because it is this sequence which is actually shared. So the RD between sampling two different finite sequences of weighted vectors from two different inputs is an accurate assessment of how easily an observer can distinguish between samples from one text’s embedding and samples from the other text’s embedding, providing a measure of privacy.
@@ 3.3 Rényi Divergence between two Sampling Distributions
Since the theory of NVIB assumes that the posterior distribution which embeds the information about the input data is a Dirichlet Process, $Q=DP(\alpha^{q}_{0},G^{q}_{0})$, we require a sampling procedure that generates finite sequences of weighted vectors which approximate sampling from this process. For $DP(\alpha^{q}_{0},G^{q}_{0})$, defined in Equation 4, weights can be sampled using a stick-breaking process parameterized by $\alpha^{q}_{0}$, and vectors can be sampled independently by first sampling a component $i$ from the mixture distribution $G^{q}_{0}$ and then sampling a vector from that component’s distribution $G^{q}_{i}=\mathcal{N}(\bm{\mu}^{q}_{i},(\bm{\sigma}^{q}_{i})^{2})$. However, sampling from the discrete list of components $i$ makes it difficult for NVIB to perform backpropagation through the sampling step, so instead Henderson and Fehr (2023) propose a sampling method that first samples the total weight assigned to all the vectors from a given component $i$, and then samples weighted vectors independently from the component’s distribution $G^{q}_{i}$. They prove that the Dirichlet Process distribution in Equation 4 is equivalent to such a factorized distribution as follows:
$F$~~~ $=~~~\sum_{i=1}^{n+1}~{\rho}_{i}{F}_{i},$ $\bm{\rho}$~~~ $\sim~~~\operatorname*{Dir}({\alpha}^{q}_{1},\ldots,{\alpha}^{q}_{n+1}),$ ${F}_{i}$~~~ $=~~~DP(G_{i}^{q},\alpha_{i}^{q}).\vskip-4.30554pt$ (6)
@@
In particular, Henderson and Fehr (2023) show that sampling the total weights $\bm{\rho_{i}}$ across $i$ can be done by sampling from a Dirichlet distribution parameterized by $\bm{\alpha^{q}}$, and the distribution of weighted vectors sampled from each component is another Dirichlet process. To generate a sample with a finite number of vectors, they assume a given bound $\kappa_{i}$ on the number of vectors sampled from each component $G^{q}_{i}$, and sample their weights with a symmetric Dirichlet distribution $\operatorname*{Dir}(\frac{{\alpha}^{q}_{i}}{\kappa_{i}},\overset{\kappa_{i}}{\ldots},\frac{{\alpha}^{q}_{i}}{\kappa_{i}})$. In our experiments, one vector is sampled from each component, so $\kappa_{i}=1$.
@@
The different weighted vectors sampled from a Dirichlet Process specify a mixture distribution, which is permutation-invariant. However, when a sample is generated with our sampling procedure, the weighted vectors are output in a specific order. We could have a better approximation to a Dirichlet Process by randomly permuting the weighted vectors, but for simplicity we output them in the order of the tokens in the sentence. This simplifies the computation of the RD between two samples because it implies that sampled vectors for two different inputs can be aligned by their token position, which gives us an upper bound on the Dirichlet Process case, since the ordered list is more informative. [^3] To handle the case where the two examples have different numbers of tokens, we pad the inputs sequences so that they all have the same length, and then assume that pad tokens have parameters $\mu_{i}{=}\mathbf{0},~\sigma_{i}{=}\mathbf{1},~\alpha_{i}{=}\mathbf{0}$. Alternatively, we could have defined adjacent inputs as all having the same length, but this approach is more appropriate for the BDP measure and still gives a meaningful privacy measure. We leave better bounds on the RD between samples from Dirichlet Processes to future work. This gives us the following formula for calculating the RD between samples from two different input embeddings.
@@
$+~\sum_{i=1}^{n+1}\kappa_{i}\Biggl(\frac{\lambda}{2}\Biggl\lVert\frac{\bm{\mu}^{q}_{i}-\bm{\mu}^{q^{\prime}}_{i}}{\bm{\sigma^{\prime}}_{i}}\Biggr\rVert^{2}~+~\frac{1}{1-\lambda}\bm{1}\left(\log\frac{\bm{\sigma^{\prime}}_{i}}{(\bm{\sigma}_{0}^{p})^{(1-\lambda)}(\bm{\sigma}^{q}_{i})^{\lambda}}\right)\Biggr),\vskip-2.15277pt$
@@ 4 Experiments
We empirically evaluate NVDP on NLP tasks from the GLUE benchmark (Wang et al., 2018) in terms of both accuracy and privacy. This is a commonly used benchmark for evaluating performance on a variety of natural language understanding tasks.
@@ 4.1 Experimental Setup
We evaluate the performance of NVDP on the GLUE (Wang et al., 2018) benchmark, which is a collection of different natural language understanding tasks including similarity and paraphrasing tasks, text classification, and natural language inference. For natural language inference, we experiment on QNLI (Rajpurkar et al., 2016) and RTE (Dagan et al., 2005). For paraphrase detection, we evaluate on MRPC (Dolan and Brockett, 2005), STS-B (Cer et al., 2017), and QQP (Chen et al., 2018). For text classification, we evaluate on SST-2 (Socher et al., 2013).
@@
We use BERTBase, which is $12$ layers and 110110M parameters. We use the following hyper-parameters for fine-tuning BERT: a sequence length of $512$, a train batch size of $64$ and evaluation batch size of $8$. We use the stable variant of the Adam optimizer (Zhang et al., 2020; Mosbach et al., 2020) with the learning rate of 2e72e{-}7 through all experiments. We use a warm-up step of $0.2$.
@@
  • Baseline (Base): we use vanilla BERTBase model without any regularization technique.
@@
We consider several simplified versions of our proposed model and other privacy-preserving baselines to comprehensively evaluate the advantages of NVDP. These ablations and baselines allow us to isolate the impact of NVIB’s nonparametric regularization and task-aware noise calibration. One ablation (VTDP) removes the use of Bayesian nonparametrics for regularization, and two ablations provide reference models which can only share single-vector embeddings instead of multivector transformer embeddings.
  • Variational Transformer Differential Privacy (VTDP): We consider a simplified version of our proposed model where the noise introduced with NVIB is replaced with noise introduced by vector-space VIB applied to each token vector independently, which we call the VTDP model. This provides a strong reference model which is the same as our NVDP model but without the holistic regularization of the entire sequence of vectors provided by NVIB. We utilize BERTBase with an additional VIB layer to implement the information bottleneck principle outlined by Mahabadi et al. (2021). The VIB layer encodes input representations in a compressed latent space, leveraging VIB to learn a stochastic latent variable. This variable captures task-relevant information while filtering out irrelevant features, enhancing generalization. This integration enables effective fine-tuning, particularly in low-resource scenarios, by prioritizing essential information and mitigating overfitting risks. The compressed latent representation is compared to a Gaussian prior, ensuring a structured and regularized latent space. The RDP guarantee, defined between Gaussian distributions, follows the formulation in Equation 10 of Van Erven and Harremos (2014):
@@
    $=$ $\frac{\lambda}{2}\Biggl\lVert\frac{\bm{\mu}^{q}_{i}-\bm{\mu}^{q^{\prime}}_{0}}{\bm{\sigma^{\prime}}_{i}}\Biggr\rVert^{2}+\frac{1}{1-\lambda}\bm{1}\left(\log\frac{\bm{\sigma^{\prime}}_{i}}{(\bm{\sigma}_{0}^{q^{\prime}})^{(1-\lambda)}(\bm{\sigma}^{q}_{i})^{\lambda}}\right).$
@@
    where the noise scale $\sigma$ is fixed and shared across all dimensions and inputs (we use $\sigma=0.55$). In contrast to our NVDP model, which learns a data-adaptive and task-aware noise structure via NVIB, VIB-fixed applies a uniform and input-independent perturbation to a single vector. This serves as a transparent benchmark closely aligned with classical Gaussian DP and admits closed-form RD computations. For two adjacent inputs $x$ and $x^{\prime}$, the order $\lambda$ RD between their noisy distributions is given by Mironov (2017):
@@
    This provides an analytically grounded reference point for comparison with NVDP, allowing us to report summary statistics of these pairwise RD across all evaluated embeddings.
@@
    Thus, VIB-learned samples from a Gaussian distribution with mean $\mu(x)$ and diagonal covariance $\mathrm{diag}(\bm{\sigma}^{2})$, where the noise magnitudes are learned globally during training but remain input-independent. This parallels the VIB-fixed mechanism while enabling dimension-wise adaptive perturbation. For two adjacent inputs $x$ and $x^{\prime}$, the order $\lambda$ RD between their privatized distributions is (Mironov, 2017):
    $D_{\lambda}\!\left(\mathcal{N}(\mu(x),\mathrm{diag}(\bm{\sigma}^{2}))\,\big\|\,\mathcal{N}(\mu(x^{\prime}),\mathrm{diag}(\bm{\sigma}^{2}))\right)=\frac{\lambda}{2}\left\lVert\frac{\mu(x)-\mu(x^{\prime})}{\bm{\sigma}}\right\rVert_{2}^{2}.$ (10)